OpenStudy (ender):

Simplify (4^7 over 5^2)^3

OpenStudy (dallascowboys88):

im helping him already sorry bro.he asked me

OpenStudy (ender):

I'm a girl

OpenStudy (dallascowboys88):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{(a^b \times c^d)^e=a^{be} \times c^{de}}}\]

OpenStudy (dallascowboys88):

im helping her already she asked me first.im already hellping @Ender

OpenStudy (jiteshmeghwal9):

okay go on

OpenStudy (dallascowboys88):

thanks

OpenStudy (jiteshmeghwal9):

I didn't notice her tagging u

OpenStudy (dallascowboys88):

@jiteshmeghwal9

OpenStudy (dallascowboys88):

its okay

OpenStudy (jiteshmeghwal9):

...

OpenStudy (dallascowboys88):

the options?

OpenStudy (ender):

A. 4^10 Over 5^5 B. 4^4 Over 5 C. 4^21 over 5^6 D. 12^7 over 15^2

OpenStudy (ender):

I think it's C

OpenStudy (dallascowboys88):

ur right.see ur good at this.

OpenStudy (ender):

thanks

OpenStudy (dallascowboys88):

np medal?

OpenStudy (dallascowboys88):

give medal.plz

OpenStudy (dallascowboys88):

lol i answered her question,but no medal from here thats weird. :/